OpenStudy (saylilbaby):

what are the shifts for the graph of (x)=1/x+5-4?

OpenStudy (saylilbaby):

@Astrophysics @TheSmartOne

OpenStudy (astrophysics):

Do you mean \[f(x) = \frac{ 1 }{ x }+1\]

OpenStudy (astrophysics):

Your question isn't clear to me

TheSmartOne (thesmartone):

or maybe: \[f(x) = \frac{ 1 }{ x+5 }-1\]

OpenStudy (astrophysics):

OOOOOH

OpenStudy (astrophysics):

Well the (x+5) would tell you it's a horizontal shift in the left axis, and the -1 would be vertical units, can you figure out if it's up or down? Note the graph of f(x) = 1/x looks as such  it's asymptotic